OpenStudy (anonymous):

2 4 - + - 9 5 Write As A Mixed Number

OpenStudy (anonymous):

@jazy

OpenStudy (anonymous):

What is the LCD(Least common denominator)?

OpenStudy (anonymous):

45

OpenStudy (anonymous):

Right! So, what are the numerators? 2/9 = ?/45 (You multiplied by 5) 4/5 = ?/45 (You multiplied by 9)

OpenStudy (anonymous):

2/9=0.2222 Repeatedly 4/5=0.8

OpenStudy (anonymous):

No. Sorry here is what I meant: \[\frac{ 2 }{ 9 } \to \frac{ ? }{ 45 }\]\[\frac{ 4 }{ 5 } \to \frac{ ? }{ 45 }\]

OpenStudy (anonymous):

So Do I Cross Multiply

OpenStudy (anonymous):

@jazy

OpenStudy (anonymous):

That could work too. Or you can think For the fraction 2/9.... 9 * 5 = 45 so multiply 2 * 5 = ? For the fraction 4/5.... 5 * 9 = 45 so multiply 4 * 9 = ?

OpenStudy (anonymous):

i Got 10 And 36

OpenStudy (anonymous):

@jazy

OpenStudy (anonymous):

Right, so add the fractions: \[\frac{ 10 }{ 45 } + \frac{ 36 }{ 45 }\] What fraction are you left with?

OpenStudy (anonymous):

46 - 45

OpenStudy (anonymous):

Ok so how many times does 46 go into 45? How much is left over?

OpenStudy (anonymous):

None

OpenStudy (anonymous):

Sorry I meant, how many times does 45 go into 46? Now how much is left over?

OpenStudy (anonymous):

it Goes In There 1 Time

OpenStudy (anonymous):

Right so your whole number will be 1. How many is left over?

OpenStudy (anonymous):

44 - 45

OpenStudy (anonymous):

No. What is 46 - 45?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Okay it goes in ONE time and has 1 left over. Remember your denominator was 45... so: \[1\frac{ 1 }{ 45 }\]

OpenStudy (anonymous):

i Got It Thannks ,

OpenStudy (anonymous):

Welcome(: